Explain how to solve equations involving absolute terms by solving following problem:
| x - ¼ | = | ½ x - ¾ |


Answer by stanbon(75887)

Explain how to solve equations involving absolute terms by solving following problem:
| x - ¼ | = | ½ x - ¾ |
-----------------------------
Assume x-(1/4) and (1/2)x-(3/4) are both positive
Then you get:
x-(1/4) = (1/2)x-(3/4)
(1/2)x = (-2/4)
x = -1
---------------
Assume x-(1/4) is negative and (1/2)x-(3/4) is positive:
Then you get:
-x+(1/4) = (1/2)x-(3/4)
(3/2)x = 1
x = 2/3
Testing x = 2/3 you get:
|2/3 - 1/4| = |(1/2)(2/3)-(3/4)|
|(5/12)| = |(1/3)-(3/4)|
5/12 = |-5/12|
5/12 = 5/12
True
----------------
Final solution:
x = 2/3 or x = -1
======================
Cheers,
Stan H.